Iucifer:

Type the correct answer in the box. Use numerals instead of words. If necessary, use / for the fraction bar. Find the remainder of (h4 + h2 – 2) ÷ (h + 3). The remainder is

Angee:

I can't tell you the answer but i suggest take a way the h's them do what is asks and pout it back on at the end like (4/2-2)+3 then once you got that add the 3

jhonyy9:

so just divide (h^4 +h^2 -2) ÷ (h +3) = ?

Extrinix:

@surjithayer wrote:
h+3=0 h=-3 -3)1 0 1 0 -2( | -3 9 -30 +90 | 1 -3 10 -30 |88 \[Quotient=x^3-3x^2+10x-30\] Remainder=88
explanation?

surjithayer:

h+3)h^4+0h^3+h^2+0h-2(h^3-3h^2+10h-30 h^4+3h^3 - - ___________________ -3h^3+h^2 -3h^3-9h^2 + + _____________________ 10h^2+0h 10h^2+30h - - ________________________ -30h-2 -30h-90 + + __________ 88 ___________ Earlier i have solved by synthetic division.

surjithayer:

h+3=0 h=-3 -3)1 0 1 0 -2( | -3 9 -30 +90 | 1 -3 10 -30 |88 \[Quotient=h^3-3h^2+10h-30\] Remainder=88